# Capturing the Elusive Curve-Crossing in Low-Lying States of Butadiene With Dressed TDDFT

### v2

@@ I Introduction
Linear polyenes have long been the focus of intensive experimental and theoretical work owing to their importance in biological and technologically important systems Hudson and Kohler (1972); Schulten and Karplus (1972); Schulten et al. (1976); Andrews and Hudson (1978); Cave and Davidson (1987); Orlandi et al. (1991); Fuss et al. (1998); Christensen et al. (2004); Levine and Martínez (2007); Levine and MartÃnez (2009); Komainda et al. (2013); Barford (2022). In particular the photophysical behavior of the smallest polyenes, s-trans-butadiene and s-trans-hexatriene, is of significant interest in organic electronics and photovoltaics Papagiannakis et al. (2002), and are viewed as representative motifs of polyenes responsible for triggering vision Orlandi et al. (1991). Despite this longstanding interest, accurately describing the electronic structure and photodynamics of linear polyenes has proven challenging for theoretical methods, leading to unreliable predictions of relaxation processes through conical intersections after photo-excitation. These difficulties stem from the nature of the low-lying excited states in these molecules, particularly the dark 21Ag state, due to its partial doubly-excited character Ohmine et al. (1978); Lappe and Cave (2000), which many single reference methods struggle to capture. Conventional computational techniques such as complete active space self-consistent field (CASSCF) and equation of motion coupled cluster with singles and doubles (EOM-CCSD) frequently fall short in describing these excited states Loos et al. (2019); Park et al. (2021). Even though these approximate methods can be adjusted to describe a single-point calculation of 21Ag accurately, their reliability across geometry variations is suspect. They typically incorrectly predict the crossing of the potential energy surfaces of the two lowest excited states, 11Bu and 21Ag Park et al. (2021). As a result, many studies have struggled to describe the ultrafast dynamics of polyenes with the precision required for predicting observed results.
@@
In principle, time-dependent density functional theory (TDDFT) Runge and Gross (1984); Casida (1995); Petersilka et al. (1996) could serve as a low-cost alternative due to its famously tolerable scaling. However, in practice the adiabatic approximation inherent in its standard linear response formulation ignores double excitation contributions, and as a result it struggles to accurately describe the energy ordering and shapes of the surfaces for these states Park et al. (2021). Going beyond adiabatic TDDFT (ATDDFT), so-called dressed TDDFT approaches have been developed Tozer et al. (1999); Tozer and Handy (2000); Cave et al. (2004); Casida (2005); Mazur and Włodarczyk (2009); Mazur et al. (2011); Romaniello et al. (2009); Sangalli et al. (2011); Huix-Rotllant et al. (2011); Casida and Huix-Rotllant (2012, 2016); Maitra (2022); Dar and Maitra (2023), in which a frequency-dependent kernel is designed to incorporate contributions from double excitations. The original version operated within the Tamm-Dancoff approximation Maitra et al. (2004); Cave et al. (2004); Casida (2005); Mazur and Włodarczyk (2009); Casida and Huix-Rotllant (2016), denoted here as dressed Tamm-Dancoff (DTDA), and has successfully predicted the energies of double excitations across a variety of molecules Cave et al. (2004); Huix-Rotllant et al. (2011); Mazur and Włodarczyk (2009); Mazur et al. (2011); Maitra (2022). However, because the Tamm-Dancoff approximation inherently does not preserve oscillator strengths, DTDA cannot be used reliably for transition properties including the non-adiabatic couplings between states needed for dynamics. In Ref. Dar and Maitra (2023), we derived a frequency-dependent kernel designed to function within the full TDDFT linear response framework, ensuring that the oscillator strength sum rule is preserved. Its performance on model as well as real systems, including the lowest ${}^{1}D$ excitations of the Be atom and LiH molecule across varying interatomic distances was found to be satisfactory. This kernel redistributes the transition density of a Kohn-Sham (KS) single excitation into a mix of single and double excitations, resulting in better predictions for transitions between excited states, as demonstrated in Ref. Dar and Maitra (2023). The examples in Ref. Dar and Maitra (2023) involved cases where one single excitation couples with a double excitation while more commonly, and in the case of the $2^{1}$g state of linear polyenes, the state is composed of several single-excitations mixing with a double-excitation with respect to the ground-state KS reference. Here we extend the approach of Ref Dar and Maitra (2023) to scenarios where multiple single excitations couple with a double excitation. We show that the potential energy surfaces of 11Bu and 21Ag for s-trans-butadiene resulting from our dressed TDDFT (DTDDFT) display a crossing close to that predicted by the highly-accurate $\delta$-CR-EOMCC(2,3) Piecuch et al. (2015); Park et al. (2021) method, which we use as a benchmark in this study.
@@ II Theory
where $\nu_{Q}$ and $\nu_{D}\equiv\nu_{s1}+\nu_{s2}$ are the KS frequencies of the single and double excitations, obtained from KS orbital energy differences (e.g. $\nu_{Q}=\epsilon_{a}-\epsilon_{i}$) and $\Omega_{Q(s1,s2)}^{A}$ are the ATDDFT-corrected frequencies of the single excitation $Q$ and the two single excitations $s_{1},s_{2}$ that compose the KS double-excitation. These two variants are much easier than DSMA0 to implement computationally as they require fewer two-electron matrix elements to be computed: Apart from $H_{QD}$ one can extract all the other quantities directly from the ATDDFT calculation.
@@
$\left(\sum_{q,q^{\prime}}c_{q}^{*}c_{q^{\prime}}H_{qq^{\prime}}-H_{00})\right)^{2}\Rightarrow\nu^{2}_{q}\delta_{qq^{\prime}}+4\sqrt{\nu_{q}\nu_{q^{\prime}}}f^{\rm{adia}}_{{\scriptscriptstyle\rm HXC}\,qq^{\prime}}\,$ (13)
@@
$X^{0}_{qq^{\prime}}(\omega)=\frac{H_{qD}H_{Dq^{\prime}}}{4\sqrt{\nu_{q}\nu_{q^{\prime}}}}\left[1+\frac{(H_{qq^{\prime}}+H_{DD}-2H_{00})^{2}}{\omega^{2}-(H_{DD}-H_{00})^{2}}\right]$ (15)
We dropped the last term in the denominator of Eq. (10) for the following two reasons. First, it would give the kernel an unphysical dependence on the arbitrary sign choice for the states $q,q^{\prime}$. Second, taking the Tamm-Dancoff limit of our approximation, where “backward” transitions are neglected, it should reduce to the previously-derived DTDA of Refs. Cave et al. (2004); Mazur and Włodarczyk (2009), but this only happens if this term is dropped (see the Supporting Information).
@@ II.1 Implementation
We first outline the computational process before applying DTDDFT to a particular molecular system. As in previous dressed approaches Maitra et al. (2004); Cave et al. (2004); Casida (2005); Mazur and Włodarczyk (2009); Mazur et al. (2011); Huix-Rotllant et al. (2011); Dar and Maitra (2023), the dressings in Eq. (15),(18 and (19) operate as a posteriori correction to ATDDFT calculations. Importantly, all the necessary ingredients except for $H_{qD}$ can be obtained solely from any code with ATDDFT capabilities. In atomic orbital codes, $H_{qD}$ can also be reconstructed from the ATDDFT calculation, since the required two-electron integrals are used in the computation of the kernel matrix elements. In the following, we detail this using the quantum chemistry software, NWChem but the procedure can be carried out using any code that is capable of doing ATDDFT and outputting the quantities needed to construct the matrix in Eq. (20) (or (21)) and the chosen dressing.
@@ III Example: Butadiene
We note that the earlier DTDA of Ref. Maitra et al. (2004); Cave et al. (2004) was applied to butadiene in Ref. Cave et al. (2004); Mazur and Włodarczyk (2009); Huix-Rotllant et al. (2011); Mazur et al. (2011), also with the PBE0 functional (but in smaller basis sets). While single-point calculations were performed at the ground-state geometry and at the excited state minimum, the shape of the potential energy surfaces was not studied. Interestingly, we find that DTDA with PBE0/cc-pVTZ significantly underestimates the energy of the 2Ag state throughout this range of BLA, and fails to capture the curve-crossing with 1Bu. While this is true for any of the three variants, Fig. 1 shows just the DTDAa variant. While we do not expect DTDA to provide accurate oscillator strengths of this state, calculations on systems studied previously Dar and Maitra (2023) showed that DTDA performed similarly to DTDDFT for the energies themselves, and an analysis of why this is not the case here is left for future work.
@@
Figure 1: Top panel: 11Bu (empty marker) and 21Ag (filled marker) excitation energies along the BLA coordinate of trans-butadiene from ATDDFT( 11Bu: blue and 21Ag: green) within PBE0/cc-pVTZ, “exact” reference $\delta$-CR-EOMCC(2,3) (black), and our DTDDFTa (blue crosses) curve. Middle panel: Same as in top panel but now showing DTDAa (blue crosses) 21Ag curves. Bottom panel: Norm-square of the eigenvector corresponding to 21Ag of Eq. (21) normalized according to Eq. (22).
@@ III.1 Summary
Thus, our DTDDFT approach offers a significant advance in the low-cost computational modeling of states of double-excitation character, overcoming key challenges that both ATDDFT and many wavefunction methods suffer from, and offers a more reliable foundation for studying ultrafast dynamics in molecular systems. While our earlier work Dar and Maitra (2023) had demonstrated the fundamental idea of DTDDFT, which in turn had built upon the earlier DTDA Maitra et al. (2004); Cave et al. (2004); Mazur and Włodarczyk (2009); Mazur et al. (2011); Huix-Rotllant et al. (2011) that was unable to correctly yield oscillator strengths and transition densities, here we have shown that DTDDFT can be efficiently extended and implemented to compute excitations involving several single excitations mixing with a double-excitation as in the case of realistic molecules such as the low-lying polyenes. The computational cost is not much beyond ATDDFT, involving only two-electron calculations in a small subspace, and the implementation can be straightforwardly interfaced with standard quantum chemistry codes. This makes it a promising tool for spectra, mapping out excited state potential energy surfaces, and for computing ultrafast non-adiabatic dynamics in photo-excited molecules, an increasingly topical research area.
@@ III.3.1 Reduction of DTDDFT kernel to DTDA
Here we show that the dressed frequency-dependent kernels derived above reduce to the DTDA expression from Cave et al. (2004); Mazur and Włodarczyk (2009) when backward transitions are neglected, as in the Tamm-Dancoff approach. We start with the dressing in Eq. 18) but keeping the term, $H_{qD}H_{Dq^{\prime}}$ in the denominator from Eq. (10):
@@
from which we see that the dressing term (second term in the above equation) agrees with ours once we replace $\nu_{D}^{2}+H_{qD}H_{Dq^{\prime}}$ to $\nu_{D}^{2}$. As discussed in the main text of the paper, having $H_{qD}H_{Dq^{\prime}}$ in the denominator of the dressing makes it ill-defined by giving it an arbitrary sign-dependence. Here we see another reason to neglect it: ignoring this term makes our kernel agree with the one given in Ref. Mazur and Włodarczyk (2009); Maitra et al. (2004); Cave et al. (2004) derived from the dressed Tamm-Dancoff approach.

### v3

@@ I Introduction
Linear polyenes have long been the focus of intensive experimental and theoretical work owing to their importance in biological and technologically important systems Hudson and Kohler (1972); Schulten and Karplus (1972); Schulten et al. (1976); Andrews and Hudson (1978); Cave and Davidson (1987); Orlandi et al. (1991); Fuss et al. (1998); Christensen et al. (2004); Levine and Martínez (2007, 2009); Komainda et al. (2013); Barford (2022). In particular the photophysical behavior of the smallest polyenes, s-trans-butadiene and s-trans-hexatriene, is of significant interest in organic electronics and photovoltaics Papagiannakis et al. (2002), and are viewed as representative motifs of polyenes responsible for triggering vision Orlandi et al. (1991). Despite this longstanding interest, accurately describing the electronic structure and photodynamics of linear polyenes has proven challenging for theoretical methods, leading to unreliable predictions of relaxation processes through conical intersections after photo-excitation. These difficulties stem from the nature of the low-lying excited states in these molecules, particularly the dark 21Ag state, due to its partial doubly-excited character Ohmine et al. (1978); Lappe and Cave (2000); Kossoski et al. (2024), which many single reference methods struggle to capture. Conventional computational techniques are often unreliable for these excited state: it is well-documented that equation of motion coupled cluster with singles and doubles (EOM-CCSD) falls short Loos et al. (2019); Park et al. (2021), while complete active space self-consistent field methods (CASSCF and CASPT2) need a careful evaluation of the active space Levine and Martínez (2009); Park et al. (2021). Even though these approximate methods can be adjusted to describe a single-point calculation of 21Ag accurately, their reliability across geometry variations is suspect. They typically incorrectly predict the crossing of the potential energy surfaces of the two lowest excited states, 11Bu and 21Ag Park et al. (2021). As a result, many studies have struggled to describe the ultrafast dynamics of polyenes with the precision required for predicting observed results.
@@
In principle, time-dependent density functional theory (TDDFT) Runge and Gross (1984); Casida (1995); Petersilka et al. (1996) could serve as a low-cost alternative due to its famously tolerable scaling. However, in practice the adiabatic approximation inherent in its standard linear response formulation ignores double excitation contributions, and as a result it struggles to accurately describe the energy ordering and shapes of the surfaces for these states Park et al. (2021). Other DFT-based methods that have been explored for double-excitations include orbital-optimized $\Delta$SCF Hait and Head-Gordon (2021), ensemble-DFT Yang et al. (2017); Cernatic and Fromager (2024); Gould et al. (2021), and pp-RPA Yu et al. (2024). Within TDDFT, to go beyond adiabatic TDDFT (ATDDFT) so-called dressed TDDFT approaches have been developed Tozer et al. (1999); Tozer and Handy (2000); Cave et al. (2004); Casida (2005); Mazur and Włodarczyk (2009); Mazur et al. (2011); Romaniello et al. (2009); Sangalli et al. (2011); Huix-Rotllant et al. (2011); Casida and Huix-Rotllant (2012, 2016); Maitra (2022); Dar and Maitra (2023), in which a frequency-dependent kernel is designed to incorporate contributions from double excitations. The original version operated within the Tamm-Dancoff approximation Maitra et al. (2004); Cave et al. (2004); Casida (2005); Mazur and Włodarczyk (2009); Casida and Huix-Rotllant (2016), denoted here as dressed Tamm-Dancoff (DTDA), and has successfully predicted the energies of double excitations across a variety of molecules Cave et al. (2004); Huix-Rotllant et al. (2011); Mazur and Włodarczyk (2009); Mazur et al. (2011); Maitra (2022). However, because the Tamm-Dancoff approximation inherently does not preserve oscillator strengths, DTDA cannot be used reliably for transition properties including the non-adiabatic couplings between states needed for dynamics. In Ref. Dar and Maitra (2023), we derived a frequency-dependent kernel designed to function within the full TDDFT linear response framework, ensuring that the oscillator strength sum rule is preserved. Its performance on model as well as real systems, including the lowest ${}^{1}D$ excitations of the Be atom and LiH molecule across varying interatomic distances was found to be satisfactory. This kernel redistributes the transition density of a Kohn-Sham (KS) single excitation into a mix of single and double excitations, resulting in better predictions for transitions between excited states, as demonstrated in Ref. Dar and Maitra (2023). The examples in Ref. Dar and Maitra (2023) involved cases where one single excitation couples with a double excitation while more commonly, and in the case of the $2^{1}$g state of linear polyenes, the state is composed of several single-excitations mixing with a double-excitation with respect to the ground-state KS reference. Here we extend the approach of Ref. Dar and Maitra (2023) to scenarios where multiple single excitations couple with a double excitation. We show that the potential energy surfaces of 11Bu and 21Ag for s-trans-butadiene resulting from our dressed TDDFT (DTDDFT) display a crossing close to that predicted by the highly-accurate $\delta$-CR-EOMCC(2,3) Piecuch et al. (2015); Park et al. (2021) method, which we use as a benchmark in this study.
@@ II Theory
where $\nu_{Q}$ and $\nu_{D}\equiv\nu_{s1}+\nu_{s2}$ are the KS frequencies of the single and double excitations, obtained from KS orbital energy differences (e.g. $\nu_{Q}=\epsilon_{a}-\epsilon_{i}$) and $\Omega_{Q(s1,s2)}^{A}$ are the ATDDFT-corrected frequencies of the single excitation $Q$ and the two single excitations $s_{1},s_{2}$ that compose the KS double-excitation. The s subscript on DSMAS denotes the replacement of the diagonal Hamiltonian matrix elements with KS energies, while the a subscript on DSMAa indicates replacement by ATDDFT Dar and Maitra (2023). These two variants are much easier to implement computationally than DSMA0 as they require fewer two-electron matrix elements to be computed: Apart from $H_{QD}$ one can extract all the other quantities directly from the ATDDFT calculation.
@@
$\left(\sum_{q,q^{\prime}}c_{q}^{*}c_{q^{\prime}}H_{qq^{\prime}}-H_{00})\right)^{2}\Rightarrow\nu^{2}_{q}\delta_{qq^{\prime}}+4\sqrt{\nu_{q}\nu_{q^{\prime}}}f^{\rm{adia}}_{{\scriptscriptstyle\rm HXC}\,qq^{\prime}}\,.$ (13)
@@
$X^{0}_{qq^{\prime}}(\omega)=\frac{H_{qD}H_{Dq^{\prime}}}{4\sqrt{\nu_{q}\nu_{q^{\prime}}}}\left[1+\frac{(H_{qq^{\prime}}+H_{DD}-2H_{00})^{2}}{\omega^{2}-(H_{DD}-H_{00})^{2}}\right].$ (15)
We dropped the last term in the denominator of Eq. (10) for the following two reasons. First, it would give the kernel an unphysical dependence on the choice of arbitrary sign for the states $q,q^{\prime}$ (that is, if the arbitrary sign of the state $q$ was chosen flipped, all matrix elements of $\Omega_{qq^{\prime}}$ should simply flip in order to yield the same energies). Second, taking the Tamm-Dancoff limit of our approximation, where “backward” transitions are neglected, it should reduce to the previously-derived DTDA of Refs. Cave et al. (2004); Mazur and Włodarczyk (2009), but this only happens if this term is dropped (see the Supporting Information).
@@ II.1 Implementation
Before applying DTDDFT to a particular molecular system, we first outline the computational process for implementing DTDDFT. As in previous dressed approaches Maitra et al. (2004); Cave et al. (2004); Casida (2005); Mazur and Włodarczyk (2009); Mazur et al. (2011); Huix-Rotllant et al. (2011); Dar and Maitra (2023), the dressings in Eq. (15),(18 and (19) operate as a posteriori correction to ATDDFT calculations. Importantly, all the necessary ingredients except for $H_{qD}$ can be obtained solely from any code with ATDDFT capabilities. In atomic orbital codes, $H_{qD}$ can also be reconstructed from the ATDDFT calculation, since the required two-electron integrals are used in the computation of the kernel matrix elements. In the following, we detail this using the quantum chemistry software, NWChem but the procedure can be carried out using any code that is capable of doing ATDDFT and outputting the quantities needed to construct the matrix in Eq. (20) (or (21)) and the chosen dressing. A python interface with NWChem can be found at the github referenced here  [^1] https://github.com/Dawood234/DTDDFT.
@@ III Example: Butadiene
We note that the earlier DTDA of Ref. Maitra et al. (2004); Cave et al. (2004) was applied to butadiene in Ref. Cave et al. (2004); Mazur and Włodarczyk (2009); Huix-Rotllant et al. (2011); Mazur et al. (2011), also with the PBE0 functional (but in smaller basis sets). While single-point calculations were performed at the ground-state geometry and at the excited state minimum, the shape of the potential energy surfaces was not studied. Interestingly, we find that DTDA with PBE0/cc-pVTZ significantly underestimates the energy of the 2Ag state throughout this range of BLA, and fails to capture the curve-crossing with 1Bu. While this is true for any of the three variants, Fig. 1 shows just the DTDAa variant. While we do not expect DTDA to provide accurate oscillator strengths of this state, calculations on systems studied previously Dar and Maitra (2023) showed that DTDA performed similarly to DTDDFT for the energies themselves. The larger error in DTDA compared with DTDDFT could be related to the fact that the 1Bu state is predicted by PBE0/TDA about 0.5–0.9eV higher over this geometry range than PBE0/TDDFT. This error translates into an error in the denominator in the dressing of Eq. (19), leading to the error in the DTDA prediction of the 2Ag state. A closer analysis of this is left for future work.
@@
Figure 1: Top panel: 11Bu (empty marker) and 21Ag (filled marker) excitation energies along the BLA coordinate of trans-butadiene from ATDDFT( 11Bu: blue dashed and 21Ag: green dashed) within PBE0/cc-pVTZ, “exact” reference $\delta$-CR-EOMCC(2,3) (black), and our DTDDFTa (blue crosses) curve, and DTDA (orange dash-dot). Bottom panel: Norm-square of the eigenvector corresponding to 21Ag of Eq. (21) normalized according to Eq. (22).
@@ III.1 Summary
Thus, our DTDDFT approach offers a significant advance in the low-cost computational modeling of states of double-excitation character, overcoming key challenges that both ATDDFT and many wavefunction methods suffer from, and offers a more reliable foundation for studying ultrafast dynamics in molecular systems. While our earlier work Dar and Maitra (2023) had demonstrated the fundamental idea of DTDDFT, which in turn had built upon the earlier DTDA Maitra et al. (2004); Cave et al. (2004); Mazur and Włodarczyk (2009); Mazur et al. (2011); Huix-Rotllant et al. (2011) that was unable to correctly yield oscillator strengths and transition densities, here we have shown that DTDDFT can be efficiently extended and implemented to compute excitations involving several single excitations mixing with a double-excitation as in the case of realistic molecules such as the low-lying polyenes. We observe that in our example, over most of the geometries shown, the DTDDFT error for the state of double-excitation character Ag is similar to the adiabatic TDDFT error for the single-excitation Bu, when the DTDDFT is built from the same adiabatic xc functional, suggesting that DTDDFT might be able to be used with as much confidence for these states as ATDDFT is for single excitations, although further tests would be necessary. The computational cost is not much beyond ATDDFT, involving only two-electron calculations in a small subspace, and the implementation can be straightforwardly interfaced with standard quantum chemistry codes. This makes it a promising tool for spectra, mapping out excited state potential energy surfaces, and for computing ultrafast non-adiabatic dynamics in photo-excited molecules. How to compute analytic gradients with our functional for efficient calculations of mixed quantum-classical dynamics is an important direction for future research.
@@ III.3.1 Reduction of DTDDFT kernel to DTDA
Here we show that the dressed frequency-dependent kernels derived in the main text reduce to the DTDA expression from referencesCave et al. (2004); Maitra et al. (2004); Mazur and Włodarczyk (2009) when backward transitions are neglected, as in the Tamm-Dancoff approach. We start with the dressing in Eq. (18) of the main text but keeping the term, $H_{qD}H_{Dq^{\prime}}$ in the denominator from Eq. (10):
@@
from which we see that the dressing term (second term in the above equation) agrees with ours once we replace $\nu_{D}^{2}+H_{qD}H_{Dq^{\prime}}$ with $\nu_{D}^{2}$. As discussed in the main text of the paper, having $H_{qD}H_{Dq^{\prime}}$ in the denominator of the dressing makes it ill-defined by giving it an arbitrary sign-dependence. Here we see another reason to neglect it: ignoring this term makes our kernel agree with the one given in Ref. Mazur and Włodarczyk (2009); Maitra et al. (2004); Cave et al. (2004) derived from the dressed Tamm-Dancoff approach.
III.3.2 Results for PBE, PBE0 and CAM-B3LYP
Figure 3: Comparison of DTDDFTa energies based on PBE, PBE0, and CAM-B3LYP for the adiabatic functional energies (top, middle, and lowest, respectively). As in Fig 1 of the main paper, the respective ATDDFT energies for 1Bu and 2Ag are shown, along with the reference ’exact’ from $\delta$-CR-EOMCC(2,3).
Here we investigate the performance of DTDDFTa for the butadiene curve-crossing when it is built upon three different types of adiabatic functionals: a GGA (PBE), the hybrid PBE0 that was presented in the main text, and a range-separated hybrid CAM-B3LYP. The figure shows the results computed in the def2-SVP basis. We see that while predictions from all three functionals are similar for the 1Bu state, with the most accurate being CAM-B3LYP, their DTDDFTa renditions differ more for the 2Ag state, again with CAM-B3LYP being most accurate. Although the trend of the exact curve is well-captured (unlike adiabatic PBE), DTDDFTa with PBE is a significant underestimate of about 1eV, too low to display a crossing in the right region. This underestimate can be seen to be largely inherited from the underestimate of the adiabatic energy rather than the dressing per se. On the other hand the CAM-B3LYP is remarkably accurate. However, it should be noted that we were not able to confirm that the CAM-B3LYP result shown is converged with respect to basis set (as we do for PBE0 in the Fig. 2 of the main paper); the range-separated hybrid appears to yield more sensitivity to basis set in our DTDDFT. A deeper study of why this is so is beyond the scope of the present study.